I need help solving this:
A company ABC asked you to design a simple payroll program that calculates and employee's weekly gross pay, including any overtime wages. If employees work over 40 hours in a week, they will get 1.5 times of their regular hourly rate for all hours over 40.
This is the answer that I am supposed to get $1650 . Everytime I try to solve it I get 2,250 as the weekly gross pay.
Enter weekly hours worked: 50
Enter hourly rate: 30
Weekly gross pay: $ 1650

Answer by richwmiller(17219)

50-40=10
30*1.5=45
40*30=1200
10*450=450
1200+450=1650